OpenStudy (anonymous):

Find (fg)(-4) when f(x) = x + 5 and g(x) = 5x2 + 10x - 5.

jimthompson5910 (jim_thompson5910):

f(x) = x+5 f(-4) = -4+5 f(-4) = 1 --------- g(x) = 5x^2+10x-5 g(-4) = 5(-4)^2+10(-4)-5 g(-4) = 5(16)+10(-4)-5 g(-4) = 80-40-5 g(-4) = 35 ------------------------------------------- (fg)(x) = f(x)*g(x) (fg)(-4) = f(-4)*g(-4) (fg)(-4) = 1*35 (fg)(-4) = 35

OpenStudy (anonymous):

fg=(x+5)(5x2 +10x -5) fg=(5x3 +10x2 -5x +25x2 +50x -25) fg=5x3 + 35x2 + 40x -25 fg(-4) = 5(-4)^3 +35(-4)^2 +50(-4) -25 fg(-4) = 64(5) + 35(16) + 50(-4) -25